OpenStudy (anonymous):

omg this is so fraustrating, everyones ignoreing me:( please help! medals awarded: (n^2+3n+2)/(2)÷(n+1)/(2(n+2)

OpenStudy (chaise):

When you divide a fraction by a fraction, you are essentially multiplying by the recipocal of that fraction. Eg: (a/b)/(c/d)= (a/b)*(d/c) Do you need further clarification?

OpenStudy (anonymous):

i still dont get what u mean?

OpenStudy (anonymous):

(a/b)/(c/d) = (a*d)/(b*c)

OpenStudy (turingtest):

he means that\[\huge\frac{\frac ab}{ \frac cd}=\frac ab\cdot \frac cd=\frac{ac}{bd}\]so in your problem we have\[\huge{{n^2+3n+2\over2}\over{n+1\over2(n+2)}}={n^2+3n+2\over2}\cdot{2(n+2)\over n+1}\]see anything that can cancel yet?

OpenStudy (anonymous):

oh i got it, so basicall anything that is division swutches to multiplication?

OpenStudy (turingtest):

yes, but you must be sure to take the \(reciprocal\)

OpenStudy (anonymous):

my answer I got was (n+1)^2/4 is this right?

OpenStudy (anonymous):

what do you mean by take the reciprical, flip it around right?

OpenStudy (chaise):

Not quite.. I don't think that Turing Test is right. He/she hasn't taken the reciprocal in the initial algebraic working out, although in your working out he/she has.

OpenStudy (chaise):

Yes, just swap the numerator and denominator. (number on top and bottom)

OpenStudy (turingtest):

I don't see the /4 offhand... yes that's what I mean, flip it=reciprocal

OpenStudy (turingtest):

did I fail to take the reciprocal? I don't think so... where?

OpenStudy (anonymous):

is the answer (n+1)^2/4 right?

OpenStudy (turingtest):

no I'm pretty sure I flipped the denominator, unless I'm misreading the question I get just (n+2)^2... but maybe my eyes are crossed... let me check

OpenStudy (turingtest):

\[\huge{{n^2+3n+2\over2}\over{n+1\over2(n+2)}}={n^2+3n+2\over2}\cdot{2(n+2)\over n+1}\]so far we agree? or am I misreading something?

OpenStudy (anonymous):

(n^2+3n+2)/(2)÷(n+1)/(2(n+2) This is a little unclear on the the order of operations... Is it [(n^2+3n+2)/(2)]÷[(n+1)/(2(n+2)]?

OpenStudy (anonymous):

i got the answer (n+1)^2/4

OpenStudy (anonymous):

is this answer correct or incorrect?

OpenStudy (anonymous):

Turing, assuming I'm reading the OoO correctly, (n+2)^2 is the correct answer.

OpenStudy (turingtest):

@SmoothMath I assumed the parentheses implicitly... so yeah you must have failed to get the reciprocal maybe @katlin95 ?

OpenStudy (anonymous):

yea lol

OpenStudy (anonymous):

Yeah, the division symbol seemed to imply that the parenthesis should be where you put them.

OpenStudy (turingtest):

maybe if I write it like this it is more cear...\[\large{{n^2+3n+2\over2}\div{n+1\over2(n+2)}}={n^2+3n+2\over2}\cdot{2(n+2)\over n+1}\]so work it from there and you should get my answer

OpenStudy (anonymous):

can u guys help me solve this? subtract and write the answer in simpleast form: (9x)/(3x-4)-(12)/(3x-4)

OpenStudy (anonymous):

What he said. Specifically, take notice that \[\div (n+1)/(2(n+2)) becomes * (2(n+2))/(n+1)\]

OpenStudy (turingtest):

^^^what he said @katlin95 I see you have already posted that problem separately, which is good I have to go, you should wait for an answer there please thanks, later :)